# Bias Amplification: Large Language Models as Increasingly Biased Media

### v2

@@ Appendix I Mathematical Details for the Statistical Tests
Mitigation Strategies. There are three potential strategies to mitigate model collapse: (1) real data mixing, (2) training data concatenation, and (3) synthetic data pruning. The first approach is discussed in (Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024), where retaining a small proportion of real data in the training set was found to slow but not completely prevent model collapse. Seddik et al. (2024) suggests that synthetic data should be exponentially smaller than real data to effectively halt model collapse, which has been shown to work with a GPT2-type model when mixing either 50% or 80% real data. The second strategy, examined by Gerstgrasser et al. (2024), involves concatenating real data with all synthetic data from previous generations to fine-tune the current generation. They show that this method prevents model collapse in several generative models, as indicated by cross-entropy validation loss. Lastly, Feng et al. (2024); Guo et al. (2024) proposed selecting or pruning synthetic datasets before fine-tuning the next generation. In the experiment conducted by Guo et al. (2024) with Llama-7B on a news summarization task, they showed that oracle selection of synthetic data outperformed random selection in terms of ROUGE-1 scores. However, filtering noisy samples using a RoBERTa model did not yield effective results.

### v3

@@ Appendix H Literature Review of Mitigation Strategies
There are three potential strategies to mitigate model collapse: (1) real data mixing, (2) training data concatenation, and (3) synthetic data pruning. The first approach is discussed in (Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024), where retaining a small proportion of real data in the training set was found to slow but not completely prevent model collapse. Seddik et al. (2024) suggests that synthetic data should be exponentially smaller than real data to effectively halt model collapse, which has been shown to work with a GPT2-type model when mixing either 50% or 80% real data. The second strategy, examined by Gerstgrasser et al. (2024), involves concatenating real data with all synthetic data from previous generations to fine-tune the current generation. They show that this method prevents model collapse in several generative models, as indicated by cross-entropy validation loss. Lastly, Feng et al. (2024); Guo et al. (2024) proposed selecting or pruning synthetic datasets before fine-tuning the next generation. In the experiment conducted by Guo et al. (2024) with Llama-7B on a news summarization task, they showed that oracle selection of synthetic data outperformed random selection in terms of ROUGE-1 scores. However, filtering noisy samples using a RoBERTa model did not yield effective results.
Appendix I Theoretical Intuition
In this section, we offer an intuitive look at principal drivers of bias amplification. We then illustrate these ideas using Weighted Maximum Likelihood Estimation (WMLE).
I.1 The Causes of Bias Amplification
Intuitively, bias amplification arises when the direction in which the parameters need to move to reduce the loss also coincides with the direction that increases the level of bias on average for a given task, referred to as bias projection in the following discussion for convenience. To illustrate this, consider a fine-tuning process in which the pre-trained model parameters $\theta_{t}$ can be expressed as the sum of unbiased and biased components:
$\theta_{t}=\theta_{t,\text{unbiased}}+\theta_{t,\text{biased}}.$
Specifically, we assume: (1) there exists a unique bias direction, $u$, such that $\theta$ can be decomposed into $\theta_{\text{unbiased}}$, which is orthogonal to $u$, and $\theta_{\text{biased}}$, where $|\theta_{\text{biased}}\cdot u|>0$; and (2) the extent of bias in the model is measured by $|\theta_{\text{biased}}\cdot u|$. During gradient-based optimization, the update rule is:
If $c_{t}<0$, the gradient update will reinforce the biased component, leading to bias amplification, i.e. $\Delta|\theta_{\text{biased}}|>0$. This occurs because the gradient descent step moves the parameters further in the direction of the existing bias.
Another cause is sampling error, akin to statistical approximation error Shumailov et al. (2024). If the model has a pre-existing bias, it inherently assigns higher probabilities to tokens that produce biased outputs. Consequently, during synthetic data generation, unbiased tokens—and thus unbiased samples—are more likely to be lost at each resampling step with a finite sample, though this error vanishes as the sample size approaches infinity. This overrepresents biased patterns in the synthetic data, surpassing the model’s original bias and true next-token probabilities. Sampling error thus complements bias projection by further activating biased neurons in response to the skewed dataset.
By definition, bias projection is a sufficient condition for bias amplification, while sampling error serves as a complementary factor. However, sampling error is a sufficient condition for model collapse to occur with nonzero probability Shumailov et al. (2024). This distinction might explain why bias amplification can occur without model collapse.
I.2 Statistical Simulation
To simulate a controlled setting without sampling error, we consider a statistical estimation cycle using WMLE with a large sample size of each resampling step. Specifically, we generate a pre-training dataset $\mathcal{D}_{\text{pre}}$ with 100,000 samples from a Beta$(3,2)$ distribution, representing a biased pre-training dataset. Using maximum likelihood estimation (MLE), we estimate its probability density function, yielding the pre-trained model $f_{\text{pre}}$.
Figure 10 shows the estimated distributions gradually shift toward the mean of the biased pre-training dataset at $x=0.6$, becoming progressively more peaked over generations. This occurs despite further training on samples drawn from Beta(2,2)22(2,2)( 2 , 2 ) and synthetic data generated from successive models. The distortion arises because the fine-tuning process disproportionately emphasizes regions where the pre-trained distribution assigns higher probability, leading to biased learning.
Figure 10: Weighted Maximum Likelihood Estimation over 10 generations.
For comparison, Figure 11 presents the results using standard MLE without weighting. In this case, the estimated distributions remain stable across generations, accurately representing the Beta(2,2)22(2,2)( 2 , 2 ) distribution.
Figure 11: Maximum Likelihood Estimation over 10 generations.